Mr. Cooper's minivan could travel 240 miles on 12 gallons of gasoline. He replaced parts in the minivan. Now, the minivan travels 48 more miles than it was traveling at the original rate with 16 gallons of gasoline.
Which of the following is true regarding Mr. Cooper's minivan after replacing parts?

Answer by addingup(3677)

240/12 = 20 miles per gallon
16 x 20 = 320 miles
320 + 48 = 368 miles
368/16 = 23 miles per gallon
--------------------------------------
Before: 20 miles per gallon
after replacing parts: 23 miles per gallon